OpenStudy (anonymous):

percent of change 600 to 550

OpenStudy (anonymous):

600-550/600 *100= 9.09%

hero (hero):

\[\frac{600 - 550}{600} \times 100 = \frac{5000}{600} = 8.333 \%\] That would be percent decrease by the way

OpenStudy (anonymous):

rounded?

hero (hero):

You could put it in the form of a fraction: \[8 \frac{1}{3} \%\]